A wave is described by y(x,t) = 0.6 sin(5x-19t), where x is in meters, y is in centimeters and t is in seconds. How long (in s) does it take the wave to travel 4.8 m? Round your answer to 2 decimal places.
Answer #1

k = 5 /m

w = 19 rad/s

Velocity of wave

v = w/k = 19/5 = 3.8 m/s

Now

Time = distance / speed = 4.8/ 3.8

t = 1.263 seconds

========

Do comment in case any doubt, will reply for sure . Goodluck